Operating Systems and Systems Integration

A Quick Quiz On Module 4 — Solutions

Note: you may discuss these questions with each other, and may refer to online or any written material. Please each hand in one piece of paper with your name and class on it, together with your answers.

Please use your own words; do not just copy from the notes, as I prefer to be more confident that you understand what you have written.

1. Explain why you might use tar and when you might use cpio. State TWO applications of each.

2. Why might you use xargs?

3. State ONE reason each for using top, vmstat and ldd.

Nick Urbanik <nicku(at)nicku.org>  ver. 1.1